ILAW Manufacturing company produces bulbs that last a mean of 900 hours with a standard deviation of 110 hours. what is the probability that the mean lifetime of a random sample of 15 of these bulbs is less than 850 hours?



1
Expert's answer
2021-05-18T13:27:31-0400

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{850-900}{110/\sqrt{15}}=-1.76"

"P(z<-1.76)=0.5-0.4608=0.0392"